OpenStudy (anonymous):

What is the solution of 5=17x^2-200x? (A) 1 real number (B) 2 complex numbers that are conjugates (C) 2 real unequal numbers (D) ∅ (E) none of these

OpenStudy (bibby):

lazy lazy lazy http://openstudy.com/study#/updates/550b0d3ee4b0f4dae08515e9

OpenStudy (perl):

$$\Large {17x^2-200x=5\\ \\ \iff \\ 17x^2 - 200x - 5 = 0 } $$

OpenStudy (perl):

but you can look at the discriminant to find how many roots, what type of roots

OpenStudy (studygurl14):

If you've already asked this question, and you received help, please don't ask the question again in a new tab. We're not going to do all the work for you.

OpenStudy (perl):

$$\Large { \text {Given: }~ax^2+bx+c = 0\\ \\ \text{There are three cases}\\ \begin{cases} ~b^2-4ac >0 & then~ 2~real ~unequal ~roots \\ ~b^2-4ac =0 & then~ 1~real ~root~ repeated \\ b^2-4ac <0 & then~ 2~imaginary~ roots \end{cases} } $$